At the championship basketball game the winning score & losing score were consecutive even intergers. If the sum of the scores was 182 what were the scores?
I think the answer is 90 & 92 - correct?

Found 2 solutions by CharlesG2, Alan3354:
Answer by CharlesG2(834)

At the championship basketball game the winning score & losing score were consecutive even intergers. If the sum of the scores was 182 what were the scores?
I think the answer is 90 & 92 - correct?
x, x+2
x + x + 2 = 182
2x + 2 = 182
2x = 180
x = 90
so x+2 = 92
yes you are correct

Answer by Alan3354(69443)

182/2 = 91, the center between the 2 numbers
--> 90 & 92